list the fractions from largest to smallest
2/15,3/5,5/7,3/13

Answer by Mathtut(3670)

2/15=182/1365
:
3/5 =819/1365
:
5/7=975/1365
:
3/13=315/1365
:
now you can arrange thes from largest to smallest